# RoboGSim: A Real2Sim2Real Robotic Gaussian Splatting Simulator

### v1

@@ Abstract
Efficient acquisition of real-world embodied data has been increasingly critical. However, large-scale demonstrations captured by remote operation tend to take extremely high costs and fail to scale up the data size in an efficient manner. Sampling the episodes under a simulated environment is a promising way for large-scale collection while existing simulators fail to high-fidelity modeling on texture and physics. To address these limitations, we introduce the RoboGSim, a real2sim2real robotic simulator, powered by 3D Gaussian Splatting and the physics engine. RoboGSim mainly includes four parts: Gaussian Reconstructor, Digital Twins Builder, Scene Composer, and Interactive Engine. It can synthesize the simulated data with novel views, objects, trajectories, and scenes. RoboGSim also provides an online, reproducible, and safe evaluation for different manipulation policies. The real2sim and sim2real transfer experiments show a high consistency in the texture and physics. Moreover, the effectiveness of synthetic data is validated under the real-world manipulated tasks. We hope RoboGSim serves as a closed-loop simulator for fair comparison on policy learning. More information can be found on our project page https://robogsim.github.io/.
@@ 1 Introduction
    Synthesizer and Evaluator: RoboGSim can synthesize the realistic manipulated demonstrations with novel scenes, views, and objects for policy learning. It can also work as the Evaluator to perform model evaluation in a physics-consistent manner.
@@ 3.1 Overall Architecture
As shown in Fig. 1, RoboGSim mainly includes four parts: Gaussian Reconstructor, Digital Twins Builder, Scene Composer, and Interactive Engine. Given multi-view images and MDH parameters of the robotic arm, Gaussian Reconstructor (Sec. 3.2) reconstructs scenes and objects using 3DGS, segments the robotic arm, and builds an MDH kinematic drive graph structure to enable accurate motion modeling of the robotic arm. Digital Twin Builder (Sec. 3.3) involves mesh reconstruction of the scene and objects. Through layout alignment, the asset data flow can be interconnected, facilitating the subsequent evaluation in Interactive Engine. Scene Composer (Sec. 3.4) achieves the synthesis of novel objects, scenes, and views. Interactive Engine (Sec. 3.5) synthesizes novel view/scene/object images for policy learning. It can also evaluate the policy networks in a closed-loop manner. Moreover, we can collect the manipulated data in simulation using VR/Xbox equipment of the real-world.
@@ 3.2 Gaussian Reconstructor
where $\beta_{i}$ represents the twist angle, which is the rotation around the $x$-axis from the $(i-1)$-th joint to the $i$-th joint. $a_{i}$ denotes the link length, measuring the distance along the $x$-axis from $z_{i-1}$ to $z_{i}$. $d_{i}$ is the link offset, indicating the displacement along the $z$-axis from $x_{i-1}$ to $x_{i}$. $\theta_{i}$ represents the joint angle, rotation around the $z$-axis from $x_{i-1}$ to $x_{i}$.
@@ 3.3 Digital Twins Builder
3D Assets Generation: We employ two methods to generate 3D object assets. For real-world objects, we capture high-quality multi-view images of the objects using a turntable and extract matching features with GIM [32] to address issues such as lack of texture and reflections. We then integrate the COLMAP pipeline [31] to obtain the initial SFM point cloud, which is subsequently used for reconstruction by 3DGS. Moreover, for novel objects available on the web, we initially employ Wonder3D [19] to procure geometrically consistent normal vectors and textured meshes of the objects. Subsequently, we utilize the method in GaussianEditor [4] that applies the diffusion model [30] to facilitate object reconstruction in 3DGS.
@@ 4.3 RoboGSim as Synthesizer
We perform the experiments on a challenging ring-toss task (see Fig. 5), which is divided into two subtasks: picking up the ring and tossing it onto the target. For real data, 1,000 samples are collected manually. For a fair comparison, we used 1,000 synthetic samples generated by RoboGSim. During testing, each model was tested 10 times, with three attempts allowed per trial for grasping. If all three attempts failed, the trial was marked as unsuccessful.
As shown in Tab. 1, the synthesized data used for VLA learning produces an overall 40% grasping and 50% placing success rate. In comparison, the VLA model using the real data achieves 90% grasping and 70% placing success rate. It should be noted that manual collection takes a total of 40 hours while RoboGSim only requires 4 hours for synthesis. It is promising to further scale up the data size of synthesis for further performance improvements. Fig. 5 shows the visualization of some success and failure cases. Moreover, we also illustrate some more qualitative analysis for novel scene synthesis. As shown in Fig. 4, we display the results of the physical migration of the UR5 robot arm to new scenes, including a factory, a shelf, and two outdoor environments. The high-fidelity multi-view renderings demonstrate that RoboGSim enables the robot arm to operate seamlessly across diverse scenes.
Data Collected Time Grasp Suc. Place Suc.
Real 40h 90% 70%
Sim 4h 40% 50%
Table 1: Performance comparison between real robotic and synthesized data. We use the manually collected manipulated and synthesized data to train the VLA model, respectively.
@@ 5 Conclusion and Discussion
In this paper, we built a Real2Sim2Real simulator, based on 3DGS. We also introduce the digital twin system with spatial alignment to enables 3D assert flow. With novel viewpoint, object, trajectory and scene, our RoboGSim engine can generate high-fidelity synthesized data. Additionally, due to our precise spatial alignment, RoboGSim can serve as evaluator that allows real-time online policy evaluation.
Despite its great progress, the current version of RoboGSim has several limitations. It can only simulate rigid objects and the lighting for synthesized objects is not yet fully unified with the robotic arm. Moreover, generating geometrically consistent object meshes remains challenging, which is often key to completing complex manipulation tasks. In the near future, we will explore more advanced mesh extraction methods, further expand the task categories and establish the benchmarks to comprehensively evaluate the performance across diverse scenarios.

### v2

@@ Abstract
Efficient acquisition of real-world embodied data has been increasingly critical. However, large-scale demonstrations captured by remote operation tend to take extremely high costs and fail to scale up the data size in an efficient manner. Sampling the episodes under a simulated environment is a promising way for large-scale collection while existing simulators fail to high-fidelity modeling on texture and physics. To address these limitations, we introduce the RoboGSim, a real2sim2real robotic simulator, powered by 3D Gaussian Splatting and the physics engine. RoboGSim mainly includes four parts: Gaussian Reconstructor, Digital Twins Builder, Scene Composer, and Interactive Engine. It can synthesize the simulated data with novel views, objects, trajectories, and scenes. RoboGSim also provides an online, reproducible, and safe evaluation for different manipulation policies. The real2sim and sim2real transfer experiments show a high consistency in the texture and physics. We compared the test results of RoboGSim data and real robot data on both RoboGSim and real robot platforms. The experimental results show that the RoboGSim data model can achieve zero-shot performance on the real robot, with results comparable to real robot data. Additionally, in experiments with novel perspectives and novel scenes, the RoboGSim data model performed even better on the real robot than the real robot data model. This not only helps reduce the sim2real gap but also addresses the limitations of real robot data collection, such as its single-source and high cost. We hope RoboGSim serves as a closed-loop simulator for fair comparison on policy learning. More information can be found on our project page https://robogsim.github.io/.
@@ 1 Introduction
    Synthesizer and Evaluator: RoboGSim can synthesize the realistic manipulated demonstrations with novel scenes, views, and objects for policy learning. It can also work as the Evaluator to perform model evaluation in a physics-consistent manner.The experiment results show that our generated data can achieve the same performance as real robot data, which in a way solves the sim2real gap problem. At the same time, in experiments with novel scenes and novel perspectives, our generated data is more effective than real robot data, even with the 2D Aug method. Our evaluator also partially verifying the performance of the real robot model in a closed-loop.
@@ 3.1 Overall Architecture
As shown in Fig. 1, RoboGSim mainly includes four parts: Gaussian Reconstructor, Digital Twins Builder, Scene Composer, and Interactive Engine. Given multi-view images and MDH parameters of the robotic arm, Gaussian Reconstructor (Sec. 3.2) reconstructs scenes and objects using 3DGS, segments the robotic arm, and builds an MDH kinematic drive graph structure to enable accurate motion modeling of the robotic arm. Digital Twin Builder (Sec. 3.3) involves mesh reconstruction of the scene and objects. Through layout alignment, the asset data flow can be interconnected, facilitating the subsequent evaluation in Interactive Engine. Scene Composer (Sec. 3.4) achieves the synthesis of novel objects, scenes, and views. Interactive Engine (Sec. 3.5) synthesizes novel view/scene/object images for policy learning. It can also evaluate the policy networks in a closed-loop manner. Moreover, we can collect the manipulated data in simulation using VR/Xbox equipment of the real world.
@@ 3.2 Gaussian Reconstructor
[…]
@@ 3.3 Digital Twins Builder
3D Assets Generation: We employ two methods to generate 3D object assets. For real-world objects, we capture high-quality multi-view images of the objects using a turntable and extract matching features with GIM [33] to address issues such as lack of texture and reflections. We then integrate the COLMAP pipeline [32] to obtain the initial SFM point cloud, which is subsequently used for reconstruction by 3DGS. Moreover, for novel objects available on the web, we initially employ generative 3D reconstruction methods [20, 17] to procure 3D gaussians and textured meshes of the objects. Subsequently, we utilize the method in GaussianEditor [4] that applies the diffusion model [31] to facilitate object reconstruction in 3DGS.
@@ 4 Experiments
Grasp Suc. Place Suc.
TV NV(MD) NS TV NV NS
Real 100% 30% 40% 90% 0% 20%
Real+2D AUG 80% 100% 60% 80% 0% 60%
RoboGSim 100% 70% 100% 90% 0% 90%
Table 1: Performance on ring-toss task in real world. TV denotes test view, NV is novel view and NS is the novel scene. MD means minor deviation.
Method Grasp Suc. Place Suc.
Real-to-Real 100% 90%
Real-to-RoboGSim 100% 30%
Sim-to-Real 80% 0%
RoboGSim-to-Real 100% 90%
Table 2: Cross validation between real world and simulation. “Sim-to-Real” means that testing the VLA model trained with the Isaac sim data on the real world robot arm.
@@ 4.3 RoboGSim as Synthesizer
Figure 5: RoboGSim as Synthesizer: Rendering of the same training set in novel view and novel scenes.
@@
Figure 8: Exp on Real robot: The first row, GSim_sus, represents successful cases of RoboGSim data under the test view. The second row, GSim_fail, represents failure cases of RoboGSim data under the novel view. The third row, GSim_sus, represents successful cases of RoboGSim data under the novel scene. The fourth row, GSim_fail, represents failure cases of RoboGSim data under the novel scene. The fifth row, Real_sus, represents successful cases of real-world data under the test view. The sixth row, Real_fail, represents successful cases of real-world data under the novel view. The seventh row, Real_sus, represents successful cases of real-world data under the novel scene. The eighth row, Real_fail, represents failure cases of real-world data under the novel scene. The ninth row, Sim_fail, represents failure cases of Isaac Sim data under the test view.
@@
By using the real machine distribution to guide the RoboGSim distribution, we aim to improve the model’s success rate. We perform the experiments on a challenging ring-toss task (see Fig. 6), which is divided into two subtasks: picking up the ring and tossing it onto the target. The accuracy requirement for the Z-axis when picking up the ring is within 5mm. For real data, 1,000 samples are collected manually. For a fair comparison, we used 1,000 synthetic samples generated by RoboGSim. During testing, each model was tested 10 times, with three attempts allowed per trial for grasping. If all three attempts failed, the trial was marked as unsuccessful.
As shown in Tab. 1, We compared three models: one trained with real machine data, one trained with real machine data plus 2D AUG, and one trained with RoboGSim. The comparison was made in terms of test view, novel view, and novel scene. The results show that in the test view, the generated data from RoboGSim can achieve zero-shot capability, with performance comparable to the real machine data (both at 90%). In the novel scene case, RoboGSim performed much better than real machine data, reaching 90% compared to the 60% of real machine data. In the novel view experiment, RoboGSim also had less bias compared to the real machine data model. We also compared the effect of adding 2D AUG to the real machine data. After adding AUG, the performance in the test view dropped (90% $\rightarrow$ 80%), but in the novel scene, it improved (40% $\rightarrow$ 60%). However, in the novel view, the bias increased. Pure 2D AUG lacks spatial awareness, and its performance is far worse than that of RoboGSim, which has spatial intelligence. a It should be noted that manual collection takes a total of 40 hours while RoboGSim only requires 4 hours for synthesis. It is promising to further scale up the data size of synthesis for further performance improvements. Fig. 6 shows the visualization of some success and failure cases. Moreover, we also illustrate some more qualitative analysis for novel scene synthesis. As shown in Fig. 4, we display the results of the physical migration of the UR5 robot arm to new scenes, including a factory, a shelf, and two outdoor environments. The high-fidelity multi-view renderings demonstrate that RoboGSim enables the robot arm to operate seamlessly across diverse scenes.
Method $L_{1}$ PSNR ↑
3DGS 0.01381 34.19939
2DGS 0.01798 32.36417
PGSR 0.01925 31.72386
MipNeRF360 0.02618 23.51348
Table 3: 3DGS achieves better static reconstruction than 2DGS
@@ 5 Conclusion and Discussion
In this paper, we built a Real2Sim2Real simulator, based on 3DGS. We also introduce the digital twin system with spatial alignment to enables 3D assert flow. With novel viewpoint, object, trajectory and scene, our RoboGSim engine can generate high-fidelity synthesized data. Additionally, due to our precise spatial alignment, RoboGSim can serve as evaluator that allows real-time online policy evaluation. Despite its great progress, the current version of RoboGSim has several limitations. It can only simulate rigid objects and the lighting for synthesized objects is not yet fully unified with the robotic arm. Moreover, generating geometrically consistent object meshes remains challenging, which is often key to completing complex manipulation tasks. In the near future, we will explore more advanced mesh extraction methods, further expand the task categories and establish the benchmarks to comprehensively evaluate the performance across diverse scenarios.